# Machine-Learned Potentials for Solvation Modeling

### v2

@@ II.2   Force prediction
where $\hat{\mathbf{x}},\hat{\mathbf{y}},{\rm and}\leavevmode\nobreak\ \hat{\mathbf{z}}$ are unit vectors in the $x$, $y$, and $z$ directions, respectively, and $(x_{i},y_{i},z_{i})$ are the Cartesian coordinates of atom $i$. This vector gives the local slope of the PES in each spatial direction for that atom, holding all other atomic positions $\{\mathbf{r}_{j}\}_{j\neq i}$ and all nuclear charges $\{Z_{j}\}$ fixed. To enable simultaneous learning of energies and forces, these models are commonly trained using a combined loss function that balances errors in both quantities (see Eq. 13).
Figure 4: Illustration of a water molecule showing atomic position vectors, $\{{\bf r}_{i}\}$ and the Cartesian components of the force vector, ${\bf F}_{3}$, acting on a selected atom. All vectors are shown in a fixed coordinate system.
@@
This decomposition serves several important purposes. First, it enforces size-extensivity by ensuring that the total energy scales linearly with system size. Second, it promotes transferability, as the local energy mappings can be applied to systems of varying size and composition. This formulation also closely mirrors the structure of the many-body expansion (see Section II.5).
@@
To reduce computational cost, a finite cutoff radius (typically $5\pm 1$) is introduced to limit the range of interactions. Within this approximation, the force on atom $i$ is computed by considering only nearby atoms:
@@
Table 3 categorizes various MLP architectures based on whether atomic forces are computed as the sum of gradients of per-atom energy contributions, $-\sum_{j}\nabla_{{\bf r}_{i}}E_{j}^{\rm ML}$, or as the gradient of the total energy, $-\nabla_{{\bf r}_{i}}E_{\rm total}^{\rm ML}$.
While most MLPs compute forces as analytic gradients of a scalar energy function, an alternative strategy is to train models that directly learn and predict atomic force vectors without explicitly learning the energy. These models, which can further be classified as gradient-domain or force-only are particularly attractive when accurate force labels, such as those from AIMD or geometry optimization trajectories, are available. Such models are trained using a force-based loss function that penalizes the difference between predicted and reference forces (see Eq. 12). Force-only approaches are especially useful in applications such as vibrational analysis, solvation-induced perturbations, or geometry optimization in complex environments.
II.3   Invariant and conservative forces
@@
×$\nabla\times\mathbf{F}_{i}\neq{\bf 0}\quad\Rightarrow\quad\mathbf{F}_{i}\neq-\nabla_{\mathbf{r}_{i}}E,$ (28)
@@
To illustrate why conservative force fields satisfy the zero-curl relation ×$\nabla\times\mathbf{F}=\nabla\times(-\nabla V)=\mathbf{0}$, consider a scalar potential energy function $V(x,y,z)$. The corresponding force field is
@@
×$\nabla\times\mathbf{F}$ $=$ ×$-\nabla\times\nabla V$ (30)
@@
each component of the curl vanishes. Therefore, any sufficiently smooth scalar potential yields a conservative force field with ×$\nabla\times\mathbf{F}=\mathbf{0}$.
@@
II.4   $\Delta$-machine learning
@@
A main advantage of this approach is that the baseline and target methods need not use the same geometries. In practice, the baseline-level geometry is often used as input for both $E^{\text{base}}$ and $\Delta E^{\mathrm{ML}}$, even though the reference (target) energy corresponds to a higher-level optimized structure. This flexibility allows the correction model to be trained on baseline geometries alone. As a result, for new predictions, only the baseline geometry is needed—avoiding expensive high-level structure optimization. This strategy has been successfully applied in many studies to predict quantum chemical properties across chemical space with improved accuracy and reduced cost Ramakrishnan and von Lilienfeld (2017); Gupta, Chakraborty, and Ramakrishnan (2021); Tripathy et al. (2024); Unzueta, Greenwell, and Beran (2021). Figure 5 illustrates the application of $\Delta$-ML to predict NMR chemical shifts using baseline and target quantum chemical levels.
@@
Alternatively, force labels can be used to directly train the force correction
@@ III   Taxonomy, Techniques and Performances
This classification distinguishes scalar-only models, conservative energy–force models, gradient-domain force-only models, and vector-only models that bypass energy learning from one another. Simultaneously, it groups models by architectural aspects, such as descriptor-based NN s, graph message-passing networks, kernel methods, and linear models. Organizing models along these conceptual axes clarifies how their design choices impact physical consistency, transferability, and suitability for solvation modeling.
@@
Moreover, the boundaries between some categories in the principled classification shown in Figure 6 are inherently fluid. For instance, scalar-only models can be viewed as limiting cases of energy–force models when trained solely on scalar quantities. Many models originally developed for scalar-valued energy prediction can also be extended to energy–force or vector prediction tasks, depending on the availability of labels and the choice of loss functions. While such distinctions are conceptually useful, they often blur in practice, as most MLP architectures can be flexibly trained with or without force supervision. These considerations motivate a departure from the strict quadrant-style classification presented in Figure 6 in favor of a more narrative-driven taxonomy.
@@
While both descriptor-based and end-to-end NN-MLPs share a common downstream structure that maps learned atomic embeddings through a readout layer to predict atomic energies, the key distinction lies in their input representations. Descriptor-based models operate on fixed, handcrafted features, whereas end-to-end architectures learn atomic representations directly from atomic numbers and positions via trainable message-passing or convolutional layers.
@@
Table 1: Comparison of major categories of machine-learned interatomic potentials (MLPs) in atomistic modeling. Models are grouped by descriptor type (explicit vs. learned), architecture (feedforward, graph-based, kernel, or tensor-based), and learning paradigm (energy-based, force-based, or hybrid).
@@ III.1.1   Descriptor-based NN-MLPs (Desc-NN-MLPs)
Descriptor-based NN-MLPs are among the earliest and most widely used approaches for ML-based PES modeling. These models rely on the assumption that the total molecular energy can be decomposed into atomic contributions, each determined by the atom’s local chemical environment. This environment is encoded using symmetry-preserving descriptors that are invariant to translation, rotation, and permutation of like atoms, such as ACSFs Behler and Parrinello (2007); Behler (2011a), Coulomb matrix Rupp et al. (2012), and SOAP Bartók, Kondor, and Csányi (2013).
@@
The vector $\mathbf{h}_{i}$, denoted as the hidden-layer activation for atom $i$, is a learned latent representation—also referred to as an embedding of the input geometry. It captures chemically relevant features of the local environment in a continuous, trainable space.
@@
Figure 7: Schematic of a DeepNN architecture for predicting atomic energy contributions, as employed in models such as BPNN and DeepMD. Each atom is represented by a descriptor vector, which is passed through a feed-forward neural network composed of an input layer, multiple hidden layers, and an output layer that returns the atomic energies. The same network is applied to all atoms of a given element type, and the resulting atomic energies are summed to obtain the total energy of the system. In BPNN, descriptors are predefined symmetry functions based on local atomic environments, whereas in DeepMD, descriptors are learned from the relative positions of neighboring atoms within a locally defined reference frame.
Early NN-MLPs
PES of CO/Ni(111): The earliest NN-MLP developed by Blank et al. Blank et al. (1995) employed shallow feedforward networks with architecture adapted to the input dimensionality of the system: a 2–6–1 network (2 input neurons, 6 hidden neurons, and 1 output neuron) for modeling the potential energy surface (PES) of a CO molecule adsorbed on a Ni(111) surface, and a 12–8–1 network for an H2 molecule interacting with a Si(100)(2 $\times$ 1) surface. The hidden layer used a sigmoid activation function, while the output layer was linear. Inputs to the network consisted of geometric degrees of freedom. For the CO/Ni system, these are $x$ (the lateral displacement of the CO molecule on the surface), and $\theta$ (the tilt angle of the CO molecule relative to the surface normal).
PES of H2/K($2\times 2$)/Pd(100): The study by Lorenz et al. Lorenz, Groß, and Scheffler (2004) constructed a high-dimensional, continuous, and differentiable PES for the interaction of H2 with a K-covered Pd(100) surface using a feedforward NN, enabling efficient and accurate classical MD simulations. The network employed an 8–24–18–1 architecture, consisting of 8 input nodes, two hidden layers with 24 and 18 sigmoidal units respectively, and a single linear output node, totaling 685 trainable weights. Of the 659 total energies obtained from DFT calculations, 619 were used for training and 40 were reserved for testing. The root mean squared error (RMSE) for both training and test sets remained below 0.1 eV relative to the reference DFT values.
To encode surface periodicity and reduce data redundancy, the six geometric degrees of freedom were transformed into eight symmetry-adapted input functions
$x_{1}$ $=d$
$x_{2}$ $=e^{-Z_{c}/2}$
$x_{3}$ $=\sin^{2}\theta\cos 2\phi\left(\cos G_{x}X_{c}-\cos G_{y}Y_{c}\right)e^{-Z_{c}/2}$
$x_{4}$ $=\sin^{2}\theta\cos 2\phi\left(\cos 2G_{x}X_{c}-\cos 2G_{y}Y_{c}\right)e^{-Z_{c}/2}$
$x_{5}$ $=\cos^{2}\theta\,e^{-Z_{c}/2}$
$x_{6}$ $=\left(\cos G_{x}X_{c}+\cos G_{y}Y_{c}\right)e^{-Z_{c}/2}$
$x_{7}$ $=\left(\cos 2G_{x}X_{c}+\cos 2G_{y}Y_{c}\right)e^{-Z_{c}/2}$
$x_{8}$ $=\sin^{4}\theta\cos 4\phi\left(\cos 2G_{x}X_{c}+\cos 2G_{y}Y_{c}\right)e^{-Z_{c}/2}$
Here, $X_{c}$, $Y_{c}$, and $Z_{c}$ are the center-of-mass coordinates of the H2 molecule; $d$ is the H–H bond length; $\theta$ and $\phi$ are its polar and azimuthal orientation angles; and $G_{x}=G_{y}=2\pi/a$ are the surface reciprocal lattice vectors. In a separate study, similar coordinates were applied to model the PES of H2/Rh(111)Dianat, Sakong, and Groß (2005).
The NN-MLP accurately captured the surface corrugation and dissociation barrier landscape, with barrier heights ranging from 0.18 eV to 5 eV. MD simulations performed on the NN-based PES reproduced sticking probabilities in agreement with those obtained using analytical PES benchmarks. The use of symmetry-adapted coordinates enabled efficient encoding of physical constraints and reduced the number of required training points.
The studies Blank et al. (1995); Lorenz, Groß, and Scheffler (2004) discussed above demonstrated that NNs incorporating minimal surface symmetry can effectively represent high-dimensional PESs, offering a scalable and generalizable alternative to analytical fitting approaches for surface reaction dynamics. These symmetry coordinates can be viewed as precursors to modern atom-centered symmetry functions used in contemporary NN-MLPs. Ref. Behler (2011b) systematically presents the early works on NN-modeling of PES of small molecules, molecules adsorbed on surfaces, and bulk materials.
@@
These ACSFs are evaluated over neighbors $j$, $k$ within the cutoff radius. A variety of hyperparameters $\eta,R_{s},\zeta,\lambda$ are used to generate a high-dimensional descriptor vector $\mathbf{d}_{i}$. Typical ranges include $R_{s}\in[0.5,6.0]\leavevmode\nobreak\ \text{\AA}$, $\eta\in[0.001,10]\leavevmode\nobreak\ \text{\AA}^{-2}$, and $\zeta\in[1,10]$. The radial and angular channels produce tens to hundreds of features, depending on grid resolution.
@@
Unlike ACSFs, which aggregate neighbor contributions more generically, AEVs are explicitly structured to resolve radial and angular interactions by atom and bond types, resulting in a fixed-length, chemically discriminating feature vector optimized for transferability and GPU efficiency. ANI models (e.g., ANI-1, ANI-1x, ANI-2x) are trained on extensive datasets constructed from normal-mode sampled conformations of small organic molecules (e.g., from GDB-11 with up to 8 heavy atoms), using DFT energies ($\omega$B97X/6-31G($d$)) as reference. ANI-1ccx Smith et al. (2020) is an NN-MLP trained on a diverse subset of molecular geometries with coupled-cluster (CCSD(T)/CBS) reference data, designed to achieve high accuracy in quantum property prediction for organic molecules.
@@
[…]
@@
As such, they bridge the gap between hand-designed physical models and fully end-to-end learned representations, while underscoring that accurate treatment of charge transfer, polarization, and long-range electrostatics remains a persistent challenge, even in descriptor-free, end-to-end architectures.
@@ III.1.2   MP-LNN-MLPs: End-to-end NN-MLPs with local message passing
A key architectural feature of several models in this category, such as deep potential molecular dynamics (referred as DPMD or DeePMD) Zhang et al. (2018b), is the use of a local coordinate frame. For each atom, a symmetry-preserving transformation maps its neighbors into a canonical local reference frame, enabling the model to maintain rotational, translational, and permutational invariance while learning environment-sensitive features. Other models, such as HIPNN Lubbers, Smith, and Barros (2018); Chigaev et al. (2023), achieve similar invariance using hierarchical message passing schemes built on radial basis expansions and smooth cutoff functions.
These models are particularly well suited for large-scale MD simulations, where locality and linear scaling are critical. Although they typically lack explicit modeling of long-range electrostatics or directional information, such effects can be implicitly learned from data, making MP-LNN-MLPs competitive surrogates for ab initio MD in solvated or condensed-phase systems.
@@
Each atom’s environment is encoded in a local coordinate frame, using scaled relative positions of neighbors within a cutoff $R_{c}$. The input feature vectors, $\{\mathbf{x}_{i}\}$, are based on the inverse interatomic distances $\{1/r_{ij}\}$ to preserve translational, rotational, and permutational symmetries. These features are processed by a DeepNN (typically with 5 hidden layers, e.g., 240–120–60–30–10 neurons) to yield atomic contributions, $E_{i}$, to the total energy.
@@
Instead of constructing molecular graphs, HIPNN defines interatomic interactions via smooth, distance-based filters. Each atom is assigned a learnable feature vector ${\bf x}_{i}$, which is iteratively refined through interaction layers. These layers aggregate information from neighboring atoms within a cutoff radius using sensitivity functions such as Gaussian in inverse distance
@@
Here, $\mathbf{x}_{i}^{(\ell_{n})}$ denotes the learned representation of atom $i$ from a selected hidden layer $\ell_{n}$, and the corresponding partial energy contribution $E_{i}^{(\ell_{n})}$ reflects interactions of increasing complexity (e.g., 2-body, 3-body, etc.). In HIPNN, only a subset of hidden layers—typically chosen at regular intervals (e.g., layers 1, 3, 5)—are used for energy prediction. These layers are manually designated to capture progressively higher-order interactions, and an independent readout head is positioned at the end of each layer. Architecturally, these readouts are placed like T-junctions, branching off from the main hidden layer stack to compute partial atomic energy contributions. This hierarchical summation enhances both accuracy and interpretability by explicitly modeling many-body interactions at multiple levels of resolution. The architecture of HIPNN is compared with that of a DeepNN in Figure 8.

### v3

@@ II.3  Force prediction: Direct and energy-based
where $\hat{\mathbf{x}},\hat{\mathbf{y}},{\rm and}\penalty 10000\ \hat{\mathbf{z}}$ are unit vectors in the $x$, $y$, and $z$ directions, respectively, and $(x_{i},y_{i},z_{i})$ are the Cartesian coordinates of atom $i$. This vector gives the local slope of the PES in each spatial direction for that atom, holding all other atomic positions $\{\mathbf{r}_{j}\}_{j\neq i}$ and all nuclear charges $\{Z_{j}\}$ fixed. To enable simultaneous learning of energies and forces, these models are commonly trained using a combined loss function that balances errors in both quantities (see Eq. 13).
Figure 4: Illustration of a water molecule showing atomic position vectors, $\{{\bf r}_{i}\}$, and the Cartesian components of the force vector, ${\bf F}_{3}$, acting on a selected atom. All vectors are shown in a fixed coordinate system.
@@
This decomposition serves several important purposes. First, it enforces size-extensivity by ensuring that the total energy scales linearly with system size. Second, it promotes transferability, as the local energy mappings can be applied to systems of varying size and composition. This formulation also closely mirrors the structure of the many-body expansion (see Section II.2). While the locality of atomic energy mappings (Eq. 18) ensures size extensivity and offers the possibility of transferability across systems of varying size and composition, achieving practical transferability ultimately depends on the diversity and representativeness of the training data. As highlighted in recent studies Goodwin et al. (2024), even models based on local or equivariant architectures require carefully curated datasets to generalize reliably to unseen chemical compositions or mixtures.
@@
To reduce computational cost, a finite cutoff radius is introduced to limit the range of interactions. In practice, cutoff distances between 5 and 10 Å are commonly employed, as this range has been shown to capture most intermolecular interactions effectively. Kovács et al. (2023) However, the optimal cutoff value can vary depending on the molecular system and the underlying architecture. Within this approximation, the force on atom $i$ is computed by considering only nearby atoms:
@@
Table 3 categorizes various MLP architectures based on whether atomic forces are computed as the sum of gradients of per-atom energy contributions, $\sum_{j}\nabla_{{\bf r}_{i}}E_{j}^{\rm ML}$, or as the gradient of the total energy, $\nabla_{{\bf r}_{i}}E_{\rm total}^{\rm ML}$.
While most MLPs compute forces as analytic gradients of a scalar energy function, an alternative strategy is to train models that directly learn and predict atomic force vectors without explicitly learning the energy. These models, which can further be classified as gradient-domain or force-only are particularly attractive when accurate force labels, such as those from AIMD or geometry optimization trajectories, are available. Such models are trained using a force-based loss function that penalizes the difference between predicted and reference forces (see Eq. 12). Force-only approaches are instrumental in applications such as vibrational analysis, solvation-induced perturbations, or geometry optimization in complex environments.
II.4  Invariant and conservative forces
@@
$\nabla\times\mathbf{F}_{i}\neq{\bf 0}\quad\Rightarrow\quad\mathbf{F}_{i}\neq-\nabla_{\mathbf{r}_{i}}E,$ (30)
@@
To illustrate why conservative force fields satisfy the zero-curl relation $\nabla\times\mathbf{F}=\nabla\times(-\nabla V)=\mathbf{0}$, consider a scalar potential energy function $V(x,y,z)$. The corresponding force field is
@@
$\nabla\times\mathbf{F}$ $=$ $-\nabla\times\nabla V$ (32)
@@
each component of the curl vanishes. Therefore, any sufficiently smooth scalar potential yields a conservative force field with $\nabla\times\mathbf{F}=\mathbf{0}$.
@@
II.5  $\Delta$-machine learning
@@
A main advantage of this approach is that the baseline and target methods need not use the same geometries. In practice, the baseline-level geometry is often used as input for both $E^{\text{base}}$ and $\Delta E^{\mathrm{ML}}$, even though the reference (target) energy corresponds to a higher-level optimized structure. This flexibility allows the correction model to be trained on baseline geometries alone. As a result, for new predictions, only the baseline geometry is needed, avoiding expensive high-level structure optimization. This strategy has been successfully applied in many studies to predict quantum chemical properties across chemical space with improved accuracy and reduced cost Ramakrishnan and von Lilienfeld (2017); Gupta, Chakraborty, and Ramakrishnan (2021); Tripathy et al. (2024); Unzueta, Greenwell, and Beran (2021). Figure 5 illustrates the application of $\Delta$-ML to predict NMR chemical shifts using baseline and target quantum chemical levels.
@@
Alternatively, force labels can be used to train the force correction directly
@@ III  Taxonomy, Techniques and Performances
This classification distinguishes scalar-only models, conservative energy–force models, gradient-domain force-only models, and vector-only models, all of which bypass energy learning from one another. Simultaneously, it groups models by architectural aspects, such as descriptor-based NN s, graph message-passing networks, kernel methods, and linear models. Organizing models along these conceptual axes clarifies how their design choices impact physical consistency, transferability, and suitability for solvation modeling.
@@
Moreover, the boundaries between some categories in the principled classification shown in Figure 6 are inherently fluid. For instance, scalar-only models can be viewed as limiting cases of energy–force models when trained solely on scalar quantities. Many models originally developed for scalar-valued energy prediction can also be extended to energy–force or vector prediction tasks, depending on the availability of labels and the choice of loss functions. While such distinctions are conceptually sound, they often blur in practice, as most MLP architectures can be flexibly trained with or without force supervision. These considerations motivate a departure from the strict quadrant-style classification presented in Figure 6 in favor of a more narrative-driven taxonomy.
@@
While both descriptor-based and end-to-end NN-MLPs share a standard downstream structure that maps learned atomic embeddings through a readout layer to predict atomic energies, the key distinction lies in their input representations. Descriptor-based models operate on fixed, handcrafted features, whereas end-to-end architectures learn atomic representations directly from atomic numbers and positions via trainable message-passing or convolutional layers.
@@
Table 1: Comparison of major categories of machine-learned interatomic potentials (MLIPs) in atomistic modeling. Models are grouped by descriptor type (explicit vs. learned), architecture (feedforward, graph-based, kernel, or tensor-based), and learning paradigm (energy-based, force-based, or hybrid).
@@ III.1.1  Descriptor-based NN-MLPs (Desc-NN-MLPs)
Descriptor-based NN-MLPs are among the earliest and most widely used approaches for ML-based PES modeling. These models rely on the assumption that the total molecular energy can be decomposed into atomic contributions, each determined by the atom’s local chemical environment. This environment is encoded using symmetry-preserving descriptors that are invariant to translation, rotation, and permutation of like atoms, such as ACSFs Behler and Parrinello (2007); Behler (2011), Coulomb matrix Rupp et al. (2012), and SOAP Bartók, Kondor, and Csányi (2013).
@@
The vector $\mathbf{h}_{i}$, denoted as the hidden-layer activation for atom $i$, is a learned latent representation, also referred to as an embedding of the input geometry. It captures chemically relevant features of the local environment in a continuous, trainable space.
@@
Figure 7: Schematic of a DeepNN architecture for predicting atomic energy contributions, as employed in models such as BPNN and DeepMD. Each atom is represented by a descriptor vector, which is passed through a feedforward neural network composed of an input layer, multiple hidden layers, and an output layer that returns the atomic energies. The same network is applied to all atoms of a given element type, and the resulting atomic energies are summed to obtain the total energy of the system. In BPNN, descriptors are predefined symmetry functions based on local atomic environments, whereas in DeepMD, descriptors are learned from the relative positions of neighboring atoms within a locally defined reference frame.
@@
These ACSFs are evaluated over neighbors $j$, $k$ within the cutoff radius. A variety of hyperparameters $\eta,R_{s},\zeta,\lambda$ are used to generate a high-dimensional descriptor vector $\mathbf{d}_{i}$. Typical ranges include $R_{s}\in[0.5,6.0]\penalty 10000\ \text{\AA }$, $\eta\in[0.001,10]\penalty 10000\ \text{\AA }^{-2}$, and $\zeta\in[1,10]$. The radial and angular channels produce tens to hundreds of features, depending on grid resolution.
@@
[…]
@@
$U_{\text{tot}}=\sum_{i}\left(E_{i}^{0}+\chi_{i}q_{i}+\frac{1}{2}J_{ii}q_{i}^{2}\right)+\frac{1}{2}\iint\frac{\rho(\mathbf{r})\rho(\mathbf{r^{\prime}})}{|\mathbf{r}-\mathbf{r^{\prime}}|}d\mathbf{r}d\mathbf{r^{\prime}},$ (50)
@@
As such, they bridge the gap between hand-designed physical models and fully end-to-end learned representations, while underscoring accurate treatment of charge transfer, polarization, and long-range electrostatics that remain a persistent challenge, even in descriptor-free, end-to-end architectures.
@@ III.1.2  MP-LNN-MLPs: End-to-end NN-MLPs with local message passing
A key architectural feature of several models in this category, such as deep potential molecular dynamics (referred to as DPMD or DeePMD) Zhang et al. (2018b), is the use of a local coordinate frame. For each atom, a symmetry-preserving transformation maps its neighbors into a canonical local reference frame, enabling the model to maintain rotational, translational, and permutational invariance while learning environment-sensitive features. Other models, such as HIPNN Lubbers, Smith, and Barros (2018); Chigaev et al. (2023), achieve similar invariance using hierarchical message passing schemes built on radial basis expansions and smooth cutoff functions.
These models are particularly well-suited for large-scale MD simulations, where locality and linear scaling are critical. Although they typically lack explicit modeling of long-range electrostatics or directional information, such effects can be implicitly learned from data, making MP-LNN-MLPs competitive surrogates for ab initio MD in solvated or condensed-phase systems.
@@
Each atom’s environment is encoded in a local coordinate frame, using scaled relative positions of neighbors within a cutoff $R_{c}$. The input feature vectors, $\{\mathbf{x}_{i}\}$, are based on the inverse interatomic distances $\{1/r_{ij}\}$ to preserve translational, rotational, and permutational symmetries. These features are processed by a DeepNN (typically with five hidden layers, e.g., 240–120–60–30–10 neurons) to yield atomic contributions, $E_{i}$, to the total energy.
@@
Instead of constructing molecular graphs, HIPNN defines interatomic interactions via smooth, distance-based filters. Each atom is assigned a learnable feature vector ${\bf x}_{i}$, which is iteratively refined through interaction layers. These layers aggregate information from neighboring atoms within a cutoff radius using sensitivity functions such as a Gaussian in inverse distance
@@
Here, $\mathbf{x}_{i}^{(\ell_{n})}$ denotes the learned representation of atom $i$ from a selected hidden layer $\ell_{n}$, and the corresponding partial energy contribution $E_{i}^{(\ell_{n})}$ reflects interactions of increasing complexity (e.g., 2-body, 3-body, etc.). In HIPNN, only a subset of hidden layers, typically chosen at regular intervals (e.g., layers 1, 3, 5), are used for energy prediction. These layers are manually designated to capture progressively higher-order interactions, and an independent readout head is positioned at the end of each layer. Architecturally, these readouts are placed like T-junctions, branching off from the main hidden layer stack to compute partial atomic energy contributions. This hierarchical summation enhances both accuracy and interpretability by explicitly modeling many-body interactions at multiple levels of resolution. The architecture of HIPNN is compared with that of a DeepNN in Figure 8.